Solve By completing the Square
x²-6x=-9

Answer by richwmiller(17219)

This one is funny.
x^2-6x=-9 equals x^2-6x+9=0 right?
You will see the funny part when we are done.
x^2-6x=-9
take half of 6
3
square it
9
add 9 to both sides
x^2-6x+9=-9+9
x^2-6x+9=0
Haha, we had that already.
All we had to do was make the equation equal to 0